Eastport Inc. was organized on June 5, 2018. It was authorized to issue 420,000 shares of $9 par common stock and 45,000 shares of 5 percent cumulative class A preferred stock. The class A stock had a stated value of $25 per share. The following stock transactions pertain to Eastport Inc.:

  1. Issued 23,000 shares of common stock for $14 per share.
  2. Issued 12,000 shares of the class A preferred stock for $30 per share.
  3. Issued 59,000 shares of common stock for $17 per share.

Required:

Prepare the stockholders’ equity section of the balance sheet immediately after these transactions have been recognized.
Answer #1
Stockholders' Equity
Preferred Stock 300000 =12000*25
Common Stock 738000 =(23000+59000)*9
Paid in capital in excess of SV, PS 60000 =12000*(30-25)
Paid in capital in excess of par, CS 587000 =23000*(14-9)+59000*(17-9)
Total paid in capital 1685000
Retained earnings 0
Total Stockholders' Equity 1685000